OpenStudy (anonymous):

Use synthetic substitution to evaluate the polynomial for the given value x = -3. P(x) = x^3 + 3x^2 + 4

myininaya (myininaya):

So have you ever done synthetic division? Like do you know how to set it up?

myininaya (myininaya):

For example pretend we have \[P(x)=x^2+1\] And we wanted to know what this polynomial equals when evaluated at x=-3 The way without synthetic division is \[P(-3)=(-3)^2+1=9+1=10\] The way with synthetic division: First make sure it is in descending order: \[P(x)=x^2+0x+1 \text{ note : you may need to put placeholders like I just did}\] So We have  The remainder is what we want to look at